“A train of mass 120,000 kg is moving with velocity of 10m/s and is acted upon by a forward force of 7000 N due to engine and frictional force of 750 N. What will be the velocity of the train after 1,000 seconds?”


1
Expert's answer
2022-03-01T11:12:05-0500

The Newton's second law says

"a=\frac{F_{net}}{m}=\frac{7000-750}{120000}=0.052\:\rm m/s^2"

The final speed

"v_f=v_i+at\\
=10+0.052*1000=62\:\rm m/s"